OpenStudy (anonymous):

3x+By=-1; (2,7) find the missing coefficient

OpenStudy (anonymous):

plug in the point and solve for B

OpenStudy (anonymous):

@rachelpeters2

OpenStudy (ghostgate):

You have your equation laid out for you: 3x + By = -1 - Your points are (2, 7), so plug them in: 3(2) + B(7) = -1 - Reword it. 6 + 7B = -1 - Now solve it like you would, what would your first step be? You would be doing what to both sides, keep in mind that it has to be the opposite of what's already going on(e.g. Addition to Subtraction, Subtraction to Addition, Multiplication to Division, Division to Multiplication.). I hope this helps in someway! Have a great day! ;) {---Ghostgate---}